1a. Find all the zeros of f(x) = X^3-X^2-6X
1b. How would I use the graphing calculator to verify the zeros?

Answer by ladyluck4772(11)

1a. In order to find the zeros of x^3-x^2-6x, we must first look at how many possible zeros there are. You can do this by looking at the term with the largest exponent. In this case, the largest exponent is 3. Therefore, there are 3 possible zeros. Now that we know this, we need to factor the function.

f(x)= x(x^2-x-6)

We can factor this again.

f(x)=x(x-3)(x+2)

Now set each part of the function equal to zero.

x=0

x-3=0
x=3

x+2=0
x=-2

Therefore, the three zeros are 0, 3, and -2.

1b. In order to use the graphing calculator to verify the zeros, you would input the function under the y= key. Do not enter the f(x) part, only x^3-x^2-6x because f(x)= is the same as y=. The zeros are the points everywhere the graph crosses the x-axis.